2. (a) Use LHôpitals rule to calculate him 2 cosh(2 – 1) – 22 + 2x 3 (x - 1)4 (b) Find the Taylor polynomial P4(2) of f(0)
Answer #1

solution:- 2) a) lim 2 cost(2-1) - 734291-3 (4-14 (i fom) m - lim asinh(ur) - 2042 4 (0-1) mm Sim 2 cyth (ny) -2 12 (24) zsinPay Gay = pop (n-1)° + (n = 11+ , 60–)2+; (1-33 + (1-19 2. 14(n-1) it žq (-104 f. Pa boy = 1+ (4-1) +55 (-13t7 nos lim aegh (